If tanA ,tanB,tanC are in HP. Then a^2,b^2,c^2 are in
 

Grade:11

2 Answers

Nishant Vora IIT Patna
7 years ago
tanA ,tanB,tanC are in HP

we know thattanA = \frac{\Delta }{s(s-a)}

hence s(s-a)/ deltas(s-b)/ delta,s(s-c)/ delta are in AP

hence a,b ,c are in AP
mycroft holmes
7 years ago
@Nishant Vora: the formula is \tan \frac{A}{2} = \frac{\Delta}{s(s-a)}